### v1

@@ Discussion.
We observe that the auditing outcome ($\tilde{\epsilon}$ values for PANORAMIA and $\epsilon$ for O(1)) can be different for DP models with the same $\epsilon$ values (Table 9). We hypothesize that the auditing results may relate more to level of overfitting than the target $\epsilon$ values in trained DP models. The difference between train and test accuracies could be a possible indicator that has a stronger relationship with the auditing outcome. We also observe that O(1) shows a faster increase in $\epsilon$ for DP models with higher targeted $\epsilon$ values. We believe it depends on the actual ratio of the correct and total number of predicted samples, since O(1) considers both true positives and true negatives while PANORAMIA considers true positives only. We leave these questions for future work.

### v2

@@ Discussion.
In the image data modality, we observe that the auditing outcome ($\tilde{\epsilon}$ values for PANORAMIA and $\epsilon$ for O(1)) can be different for DP models with the same $\epsilon$ values (Table 10). We hypothesize that the auditing results may relate more to level of overfitting than the target $\epsilon$ values in trained DP models. The difference between train and test accuracies could be a possible indicator that has a stronger relationship with the auditing outcome. We also observe that O(1) shows a faster increase in $\epsilon$ for DP models with higher targeted $\epsilon$ values. We believe it depends on the actual ratio of the correct and total number of predicted samples, since O(1) considers both true positives and true negatives while PANORAMIA considers true positives only. We leave these questions for future work.
In the text data modality, the membership inference attacks (MIAs) relying on loss thresholds in privacy audits, without retraining or access to canaries, may not be sufficiently strong to effectively classify members and non-members.